OpenStudy (anonymous):

[12 T to the third/6 s to the fourth ]cubed simplify expression

OpenStudy (anonymous):

\[\left( \frac{ 12T ^{3} }{ 6s^{4} } \right)^{3}\] Is this it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do i cube just the 12 and 6 or also the s and t

OpenStudy (anonymous):

Everything gets cubed!

OpenStudy (anonymous):

I've personally arrived at the simplest form of: \[\frac{ 96T^{9} }{ 7S^{12} }\]